# Recognition of Unit Segment and Polyline Graphs is $\exists\mathbb{R}$-Complete

### v2

@@ Abstract
Given a set of objects $O$ in the plane, the corresponding intersection graph is defined as follows. A vertex is created for each object and an edge joins two vertices whenever the corresponding objects intersect. We study here the case of unit segments and polylines with exactly $k$ bends. In the recognition problem, we are given a graph and want to decide whether the graph can be represented as the intersection graph of certain geometric objects. In previous work it was shown that various recognition problems are $\exists\mathbb{R}$-complete, leaving unit segments and polylines as few remaining natural cases. We show that recognition for both families of objects is $\exists\mathbb{R}$-complete.
A graph and two representations as an intersection graph, of unit segments and polylines respectively.
@@ 1 Introduction
Many real-life problems can be mathematically described in the language of graphs. For instance, consider all the cell towers in Switzerland. We want to assign each tower a frequency such that no two towers that overlap in coverage use the same frequency. This can be seen as a graph coloring problem: Every cell tower becomes a vertex, overlap indicates an edge and a frequency assignment corresponds to a proper coloring of the vertices, see Figure 1.
@@
In many contexts, we have additional structure on the graph that may or may not help us to solve the underlying algorithmic problem. For instance, it might be that the graph arises as the intersection graph of unit disks in the plane such as in the example above. In that case, the coloring problem can be solved more efficiently [10], and there are better approximation algorithms for the clique problem [11]. This motivates a systematic study of geometric intersection graphs. Another motivation is mathematical curiosity. Simple geometric shapes are easily visualized and arguably very natural mathematical objects. Studying the structural properties of intersection graphs gives insight into those geometric shapes and their possible intersection patterns.
@@
In this work, we focus on two geometric objects; unit segments and polylines with exactly $k$ bends. Although we consider both types of geometric objects natural and well studied, to the best of our knowledge the complexity of their recognition problem was left open. We close this gap by showing that both recognition problems are $\exists\mathbb{R}$-complete.
@@ 1.1 Definition and Results
Given a finite set of geometric objects $O$, we denote by $G(O)=(V,E)$, the corresponding intersection graph. The set of vertices is the set of objects ($V=O$) and two objects are adjacent ($uv\in E$) if they intersect ($u\cap v\neq\emptyset$). We are interested in intersection graphs that come from different families of geometric objects.
@@ Results.
Unit Recognition is $\exists\mathbb{R}$-complete.
@@ 1.2 Discussion
In this section, we discuss strengths and limitations of our results from different perspectives. To supply the appropriate context, we give a comprehensive list of important geometric graph classes and the current knowledge about the complexity of their recognition problems in Table 1.
@@
string NP-complete Schaefer and Sedgwick [23, 36]
@@
Figure 2: Each box represents a different geometric intersection graph class. Those marked in green can be recognized in polynomial time. Those in blue are known to be $\exists\mathbb{R}$-complete. The ones in gray are NP-complete, and the orange ones are the new results presented in this paper.
@@
We see our main contribution in refining the hierarchy of geometric graph classes for which recognition complexity is known. Both unit segments as well as polylines with $k$ bends are natural objects that are well studied in the literature. However, the recognition of the corresponding graph classes was not studied previously. Polylines with an unbounded number of bends are equivalent to strings (it is possible to show that polylines with an unbounded number of bends are as versatile as strings with respect to the types of graphs that they can represent, since the number of intersections of any two strings can always be bounded from above [36, 37]), while polylines with 00 bends are just segments. Polylines with $k$ bends thus naturally slot in between strings and segments, and their corresponding graph class is thus also an intermediate class between the class of segment intersection graphs and string graphs, as can be seen in Figure 2. By showing that recognition for polylines with $k$ bends is $\exists\mathbb{R}$-complete for all constant $k$, we see that the switch from $\exists\mathbb{R}$-completeness (segment intersection graphs) to NP-membership (string graphs) really only happens once $k$ is infinite. Similarly, unit segment intersection graphs slot in between segment and ray intersection graphs. Intuitively, recognition of a class intermediate to two classes that are $\exists\mathbb{R}$-hard to recognize should also be $\exists\mathbb{R}$-hard, and our Theorem 1 confirms this intuition in this case.
@@
One of the consequences of $\exists\mathbb{R}$-completeness is that there are no short representations of solutions known. Some representable graphs may only be representable by objects with irrational coordinates, or by rational coordinates with nominator and denominator of size at least $2^{2^{n^{c}}}$, for some fixed $c>0$. In other words, the numbers to describe the position might need to be doubly exponentially large [27] for some graphs. For “flexible” objects like polylines, rational solutions can always be obtained by slightly perturbing the representation. For more “sturdy” objects like unit segments this may not be possible, however it is known that for example unit disks admit rational solutions despite their inflexibility [28].
@@
Given the picture of Figure 2, we wish to start unraveling a deeper story. Namely, we aim to get a better understanding of when geometric recognition problems are $\exists\mathbb{R}$-complete and when they are contained in NP. Figure 2 indicates that $\exists\mathbb{R}$-hardness comes from objects that are complicated enough to avoid a complete combinatorial characterization. Unit interval graphs, interval graphs and circle chord graphs admit such a characterization that in turn can be used to develop a polynomial-time algorithm for their recognition. On the other hand, if the geometric objects are too broad, the recognition problem is in NP. The prime example is string graphs. For string graphs, it is sufficient to know the planar graph given by the intersection pattern of the strings, which is purely combinatorial information that does not care anymore about precise coordinates. Despite the fact that there are graphs that need an exponential number of intersections, it is possible to find a polynomial-size witness [36] and thus we do not have $\exists\mathbb{R}$-hardness. We want to summarize this as: recognition problems are $\exists\mathbb{R}$-complete if the underlying family of geometric objects is at a sweet spot of neither being too simplistic nor too flexible.
@@
A classic way to make an algorithmic problem easier is to consider it only on restricted input. In the context of $\exists\mathbb{R}$ and geometric graph recognition, we would like to mention the work by Schaefer [35], where it is shown that the recognition problem of segments is already $\exists\mathbb{R}$-complete for graphs of bounded degree. We conjecture that the same is true for unit segments and polylines. However, this does not follow from our reductions since we create graphs of unbounded degree.
@@
Unit Recognition is $\exists\mathbb{R}$-complete already for graphs of bounded degree.
@@
The class $\exists\mathbb{R}$ is the family of all problems that admit a polynomial-time many-one reduction to ETR. It is known that $\textsf{NP}\subseteq\exists\mathbb{R}\subseteq\textsf{PSPACE}$ [13]. The reason that $\exists\mathbb{R}$ is an important complexity class is that a number of common problems, mainly in computational geometry, have been shown to be complete for this class. Schaefer established the current name and pointed out first that several known NP-hardness reductions actually imply $\exists\mathbb{R}$-completeness [33]. Early examples are related to recognition of geometric structures: points in the plane [29, 42], geometric linkages [1, 34], segment graphs [25, 27], unit disk graphs [21, 28], ray intersection graphs [14], and point visibility graphs [14]. In general, the complexity class is more established in the graph drawing community [16, 18, 26, 35]. Yet, it is also relevant for studying polytopes [17, 31], Nash-Equilibria [6, 8, 9, 20, 38], and matrix factorization problems [15, 39, 40, 41]. Other $\exists\mathbb{R}$-complete problems are the Art Gallery Problem [3, 44], covering polygons with convex polygons [2], geometric packing [5] and training neural networks [4, 7].
@@ 1.4 Proof Techniques
The techniques used in this paper are similar to previous work. $\exists\mathbb{R}$-membership can be established straightforwardly by constructing concrete formulae or invoking a characterization of $\exists\mathbb{R}$ using real verification algorithms [19], similar to the characterization of NP. This proof can be found in Section 3.
For $\exists\mathbb{R}$-hardness, we are in essence reducing from the Stretchability problem. In this problem, we are given a simple pseudoline arrangement as an input, and the question is whether this arrangement is stretchable. A pseudoline arrangement $\mathcal{A}$ is a set of $n$ curves that are $x$-monotone. Furthermore, any two curves intersect exactly once and no three curves meet in a single point. We assume that there exist two vertical lines on which each curve starts and ends. The problem is to determine whether there exists a combinatorially equivalent (truncated) line arrangement. See Figure 3 for an example.
For the reduction establishing $\exists\mathbb{R}$-hardness of Unit Recognition, we are given a pseudoline arrangement $\mathcal{A}$, and we construct a graph that is representable by unit segments (or polylines with $k$ bends) iff $\mathcal{A}$ is stretchable. This graph is created by enhancing $\mathcal{A}$ with more curves and taking their intersection graph. By the way these additional curves are constructed, they can always be drawn as unit segments in any line arrangement combinatorially equivalent to $\mathcal{A}$, if this line arrangement is first “squeezed” into a canonical form.
For the other direction of the equivalence we need to show that if the graph is representable by unit segments, then $\mathcal{A}$ is stretchable. To help with this proof we establish helper lemmas for both unit segments and polylines stating that cycles in a graph enforce a certain order of intersections of objects around a closed curve in any realization of that graph. The details of these helper lemmas can be found in Section 2.
Using these additional lemmas we then show that the unit segments corresponding to the vertices originally representing our pseudolines must have the same combinatorial structure as our original arrangement $\mathcal{A}$. To ensure this we add so-called probes during the construction of our graph $G$. The ideas of order-enforcing cycles and probes have already been used in different contexts [14].
The main idea to show hardness of $k$-PolyLine Recognition is to enhance the construction for Unit Recognition by an additional area where the polylines representing our pseudolines have to make $k$ bends. This then ensures that the polylines have no bends in the crucial part representing the pseudoline arrangement $\mathcal{A}$, and the same arguments as for unit segments can be used to show that this witnesses the stretchability of the original pseudoline arrangement.
@@ 2 Cycle Representations
In our reduction, we will construct a graph $G$ that contains a cycle. The cycle helps us to enforce a certain structure on any geometric representation of $G$. The arguments in this section merely use that our geometric objects are Jordan arcs whose pairwise intersections consist of a finite number of connected components, and are not specific to either unit segments or polylines.
@@
We first want to introduce some notions about realizations of induced cycles by Jordan arcs. Let $G$ be a graph, and let $C\subseteq V(G)$ be a set of vertices such that the induced graph $G[C]$ is a cycle $(c_{1},\ldots,c_{n})$ of $n\geq 4$ vertices. If we are now given a geometric representation of $G$ as an intersection graph of Jordan arcs, we can define two Jordan curves “tracing” the representation $r(C)$ sufficiently close, one on the inside and one on the outside of $r(C)$. To be precise, by sufficiently close we mean that the curves are close enough to $r(C)$, such that between the two curves no other object $o\not\in r(C)$ starts or ends, no proper crossing of two such other objects occurs, and such that every crossing of another object $o$ with one of the curves implies that $o$ also intersects $r(C)$ within a small $\epsilon$-ball centered at the intersection with the curve. We call the two curves interior boundary curve and exterior boundary curve. See Figure 4 for an illustration.
@@
Figure 4: Top row: The representation $r(C)$ by segments, polylines and strings respectively. Bottom row: The representation of the top row with two boundary curves added. One on the interior and one on the exterior of the representation. From left to right we see unit segments, polylines with one bend and general Jordan arcs. Left: The interior curve traces the Jordan arcs $abcdcdefg$. The exterior curve traces $ababcbcdedefefgfgag$. Middle: The interior curve traces the Jordan arcs $abcdcdef$. The exterior curve traces $ababcdcdefefaf$. Right: The interior curve traces the Jordan arcs $abcdcdcd$. The exterior curve traces $ababcbcdad$.
Observation 3.
The interior boundary curve lies completely within the bounded cell bounded by the exterior boundary curve.
Given either the interior or exterior boundary curve, we can record the order of traced elements, i.e., the order of elements $c\in C$ for which the boundary curve is close to $r(c)$, see Figure 4 for an illustration. This order admits some simple structure, that we describe in the following lemma.
@@
Let $C$ be an induced cycle of length $n\geq 4$, and let $b$ be either the interior or exterior boundary curve of $r(C)$. Let $s\in C^{*}$ be the string describing the cyclic order of traced elements of $b$. Then this string can be decomposed into $n$ consecutive parts $s_{1},\ldots,s_{n}$, where in each part $s_{i}$, only the elements $c_{i},c_{i+1}$ occur (note that we take indices modulo $n$).
@@
We stretch out the boundary curve horizontally and draw the Jordan arcs it traces on the top of the boundary curve (see Figure 5). Let us now assume that $c_{i}$ appears again on the sequence after $c_{i+2}$ was already seen, but not $c_{i+3}$ yet. Note that the vertices $c_{i+2}$ and $c_{i}$ are not adjacent in $C$, thus their curves are not allowed to intersect. By definition the boundary curve is also not intersected. Thus $c_{i+2}$ is enclosed by $c_{i}$ and the boundary curve and cannot intersect $c_{i+3}$. This is a contradiction, and we thus know that for all $i$, $c_{i}$ can only be seen before $c_{i+2}$ or after $c_{i+3}$ (with all indices taken modulo $n$).
We can thus split the string into $n$ consecutive parts $s_{1},\ldots,s_{n}$ by simply ending part $s_{i}$ whenever $c_{i}$ occurs for the last time. ∎
Figure 5: The Jordan arc $c_{i+2}$ is encapsulated by the boundary curve and $c_{i}$ and thus cannot intersect $c_{i+3}$.
Next, we want to argue that for our graph $G$ the realizations of certain vertices of the graph must be contained in some cell.

### v3

@@ Abstract
Given a set of objects $O$ in the plane, the corresponding intersection graph is defined as follows. Each object defines a vertex and an edge joins two vertices whenever the corresponding objects intersect. We study here the case of unit segments and polylines with exactly $k$ bends. In the recognition problem, we are given a graph and want to decide whether the graph can be represented as an intersection graph of certain geometric objects. In previous work it was shown that various recognition problems are $\exists\mathbb{R}$-complete, leaving unit segments and polylines among the few remaining natural cases where the recognition complexity remained open. We show that recognition for both families of objects is $\exists\mathbb{R}$-complete.
A graph and two representations as an intersection graph, one using unit segments and one using polylines.
@@ 1 Introduction
Many real-life problems can be mathematically described in the language of graphs. For instance, consider all the cell towers in Switzerland. We want to assign each tower a frequency such that no two towers that overlap in coverage use the same frequency. This can be seen as a graph coloring problem: Every cell tower becomes a vertex, overlap indicates an edge and a frequency assignment corresponds to a proper coloring of the vertices; see Figure˜1.
@@
In many contexts, we have additional structure on the graph that may or may not help us to solve the underlying algorithmic problem. For instance, it might be that the graph arises as an intersection graph of unit disks in the plane such as in the example above. In that case, the coloring problem can be solved more efficiently [10], and there are better approximation algorithms for the clique problem [11]. This motivates a systematic study of geometric intersection graphs. Another motivation is mathematical curiosity. Simple geometric shapes are easily visualized and arguably very natural mathematical objects. Studying the structural properties of intersection graphs gives insight into those geometric shapes and their possible intersection patterns.
@@
In this work, we focus on two geometric objects; unit segments and polylines with exactly $k$ bends. Although we consider both types of geometric objects natural and well studied, the complexity of their recognition problem was left open. We close this gap by showing that both recognition problems are $\exists\mathbb{R}$-complete.
@@ 1.1 Definition and Results
Given a finite set of geometric objects $O$, we denote by $G(O)=(V,E)$, the corresponding intersection graph. The set of vertices is the set of objects ($V=O$) and two objects $u$ and $v$ are adjacent ($uv\in E$) if they intersect ($u\cap v\neq\emptyset$). We are interested in intersection graphs that come from different families of geometric objects.
@@ Results.
Unit Segment Recognition is $\exists\mathbb{R}$-complete.
@@ 1.2 Discussion
In this section, we discuss strengths and limitations of our results from different perspectives. To supply the appropriate context, we give a comprehensive list of important geometric graph classes and the current knowledge about the complexity of their recognition problems in Table˜1.
@@
string NP-complete Kratochvíl [25], Schaefer et al. [39]
@@
segments in 3D $\exists\mathbb{R}$-complete Evans et al. [20]
Figure 2: Each box represents a different geometric intersection graph class. Upward edges between classes express containment. Those marked in green can be recognized in polynomial time. Those in blue are known to be $\exists\mathbb{R}$-complete. The ones in gray are NP-complete, and the orange ones are the new results presented in this paper.
@@
We see our main contribution in refining the hierarchy of geometric graph classes for which recognition complexity is known. Both unit segments as well as polylines with $k$ bends are natural objects that are well studied in the literature. However, the recognition of the corresponding graph classes has not been studied previously. Polylines with an unbounded number of bends are equivalent to strings (it is possible to show that polylines with an unbounded number of bends are as versatile as strings with respect to the types of graphs that they can represent, since the number of intersections of any two strings can always be bounded from above [39, 40]), while polylines with 0 bends are just segments. Polylines with $k$ bends thus naturally slot in between strings and segments, and their corresponding graph class is thus also an intermediate class between the class of segment intersection graphs and string graphs, as can be seen in Figure˜2. By showing that recognition for polylines with $k$ bends is $\exists\mathbb{R}$-complete for all constant $k$, we see that the switch from $\exists\mathbb{R}$-completeness (segment intersection graphs) to NP-membership (string graphs) really only happens once $k$ is infinite. Similarly, unit segment intersection graphs slot in between segment and ray intersection graphs. Intuitively, it is natural to expect that recognition of a class intermediate to two classes that are $\exists\mathbb{R}$-hard to recognize is $\exists\mathbb{R}$-hard again (although in general this intuition is wrong). Our Theorem˜1 confirms this intuition in the case of $k$-bend polylines with $k\in\{1,2,\dots\}\cup\{\infty\}$.
@@
One of the consequences of $\exists\mathbb{R}$-completeness is that there are no known ways to encode a solution (a representation) in polynomially many bits. Some representable graphs may only be representable by objects with irrational coordinates, or by rational coordinates with nominator and denominator of size at least $2^{2^{n^{c}}}$, for some fixed $c>0$. In other words, the numbers to describe the position might need to be doubly exponentially large [29] for some graphs. For “flexible” objects such as polylines, rational solutions can always be obtained by slightly perturbing the representation. For more “sturdy” objects such as unit segments this may not be possible, however it is known that for example unit disks admit rational solutions despite their inflexibility [30].
@@
Given the picture of Figure˜2, we aim to get a better understanding of when geometric recognition problems are $\exists\mathbb{R}$-complete and when they are contained in NP. Figure˜2 indicates that $\exists\mathbb{R}$-hardness comes from objects that are complicated enough to avoid a complete combinatorial characterization. Unit interval graphs, interval graphs and circle chord graphs admit such a characterization that in turn can be used to develop a polynomial-time algorithm for their recognition. On the other hand, if the geometric objects are too broad, the recognition problem is in NP. The prime example is string graphs. For string graphs, it is sufficient to know the planar graph given by the intersection pattern of the strings, which is purely combinatorial information that does not care anymore about precise coordinates. Despite the fact that there are graphs that need an exponential number of intersections, it is possible to find a polynomial-size witness [39] and thus we do not have $\exists\mathbb{R}$-hardness. We want to summarize this as: recognition problems are $\exists\mathbb{R}$-complete if the underlying family of geometric objects is at a sweet spot of neither being too simplistic nor too flexible.
@@
A classic way to make an algorithmic problem easier is to consider it only on a restricted input. In the context of $\exists\mathbb{R}$ and geometric graph recognition, we would like to mention the work by Schaefer [37], who showed that the recognition problem of segments is already $\exists\mathbb{R}$-complete for graphs of bounded degree. We conjecture that the same is true for unit segments and polylines. However, this does not follow from our reductions since we create graphs of unbounded degree.
@@
Unit Segment Recognition is $\exists\mathbb{R}$-complete already for graphs of bounded degree.
@@
The class $\exists\mathbb{R}$ is the family of all problems that admit a polynomial-time many-one reduction to ETR. It is known that $\textsf{NP}\subseteq\exists\mathbb{R}\subseteq\textsf{PSPACE}$ [13]. Note that there are oracle separations that indicate that $\exists\mathbb{R}$ and PSPACE are distinct [22]. The reason that $\exists\mathbb{R}$ is an important complexity class is that a number of common problems, mainly in computational geometry, have been shown to be complete for this class. The compendium by Schaefer, Cardinal and Miltzow [38] give a comprehensive overview. Schaefer established the current name and pointed out first that several known NP-hardness reductions actually imply $\exists\mathbb{R}$-completeness [35]. Early examples are related to recognition of geometric structures: points in the plane [31, 45], geometric linkages [1, 36], segment graphs [27, 29], unit disk graphs [23, 30], ray intersection graphs [14], and point visibility graphs [14]. In general, the complexity class is more established in the graph drawing community [16, 18, 28, 37]. Yet, it is also relevant for studying polytopes [17, 33], Nash equilibria [6, 8, 9, 21, 41], and matrix factorization problems [15, 42, 43, 44]. Other $\exists\mathbb{R}$-complete problems are the Art Gallery problem [3, 47], covering polygons with convex polygons [2], geometric packing [5] and training neural networks [4, 7].
@@ 1.4 Proof Techniques
The techniques used in this paper are similar to previous work. $\exists\mathbb{R}$-membership can be established straightforwardly by constructing concrete formulae or invoking a characterization of $\exists\mathbb{R}$ using real verification algorithms [19], similar to the characterization of NP. This proof can be found in Section˜3.
For $\exists\mathbb{R}$-hardness, we are in essence reducing from the Stretchability problem. In this problem, we are given a simple pseudoline arrangement as an input, and the question is whether this arrangement is stretchable. A pseudoline arrangement is a set of curves that are $x$-monotone. Furthermore, any two curves intersect exactly once and no three curves meet in a single point. We assume that there exist two vertical lines on which each curve starts and ends. The problem is to determine whether there exists a combinatorially equivalent (truncated) line arrangement. See Figure˜3 for an example.
For the reduction establishing $\exists\mathbb{R}$-hardness of Unit Segment Recognition, we are given a pseudoline arrangement $\mathcal{A}$, and we construct a graph that is representable by unit segments (or polylines with $k$ bends) if and only if $\mathcal{A}$ is stretchable. This graph is created by enhancing $\mathcal{A}$ with more curves and taking their intersection graph. Due to the way these additional curves are constructed, they can always be drawn as unit segments in any line arrangement $\mathcal{A}^{\prime}$ combinatorially equivalent to $\mathcal{A}$, if $\mathcal{A}^{\prime}$ is first “squeezed” into a canonical form.
For the other direction of the equivalence we need to show that if the graph is representable by unit segments, then $\mathcal{A}$ is stretchable. To aid with this proof, we establish helper lemmas for both unit segments and polylines stating that cycles in a graph enforce a certain order of intersections of objects around a closed curve in any realization of that graph. The details of these helper lemmas can be found in Section˜2.
Using these additional lemmas, we then show that the unit segments corresponding to the vertices originally representing our pseudolines must have the same combinatorial structure as our original arrangement $\mathcal{A}$. To ensure this, we add so-called probes during the construction of our graph $G$. The ideas of order-enforcing cycles and probes have already been used in a different context [14].
The main idea to show hardness of $k$-PolyLine Recognition is to enhance the construction for Unit Segment Recognition by an additional region where the polylines representing our pseudolines have to make their $k$ bends. This then ensures that the polylines have no bends in the crucial part representing the pseudoline arrangement $\mathcal{A}$, and the same arguments as for unit segments can be used to show that this witnesses the stretchability of the original pseudoline arrangement.
@@ 2 Cycle Representations
In our reduction, we will construct a graph $G$ that contains a cycle. The cycle helps us to enforce a certain structure on any geometric representation of $G$. The arguments in this section merely use that our geometric objects are Jordan arcs who are piecewise differentiable, and are not specific to either unit segments or polylines.
@@
We first introduce some notion about realizations of induced cycles by unit segments, and then explain how to generalize these notions to all Jordan arcs. Let $G$ be a graph, and let $C\subseteq V(G)$ be a set of vertices such that the induced graph $G[C]$ is a cycle $(c_{1},\ldots,c_{n})$ of $n\geq 4$ vertices. Consider now a given geometric representation of $G$ as an intersection graph of unit segments. We observe that the set $r(C)$ splits the plane into exactly two regions, since it is a set of unit segments whose intersection pattern is specified by the cycle $C$. We can then define two Jordan curves “tracing” the representation $r(C)$ sufficiently close, one on the inside and one on the outside of $r(C)$. By sufficiently close we mean that the curves are close enough to $r(C)$ such that (i) between the two curves no other object $o\not\in r(C)$ starts or ends, (ii) no proper crossing of two such other objects occurs, and (iii) every crossing of another object $o$ with one of the curves implies that $o$ also intersects $r(C)$ within a small $\epsilon$-ball centered at the intersection with the curve. We call the two curves interior boundary curve and exterior boundary curve. See Figure˜4 (left column) for an illustration. Given either the interior or exterior boundary curve, we can record the order of traced elements, i.e., the order of elements $c\in C$ for which the boundary curve is close to $r(c)$. We call this order the trace of the curve.
@@
Figure 4: Top row: Representations $r(C)$ by unit segments, 11-polylines and strings, respectively. Bottom row: The representation of the top row with two boundary curves added; one on the interior and one on the exterior of the representation. Left: The interior curve traces the Jordan arcs abcdcdefgabcdcdefg. The exterior curve traces ababcbcdedefefgfgagababcbcdedefefgfgag. Middle: The interior curve traces the Jordan arcs abcdcdefabcdcdef. The exterior curve traces ababcdcdefefafababcdcdefefaf. Right: The interior curve traces the Jordan arcs abcdcdcdabcdcdcd. The exterior curve traces ababcbcdadababcbcdad.
In the more general case that the geometric realization of $G$ is given in terms of Jordan curves instead of only unit segments, we can sometimes extend the same notion of interior and exterior bounding curve and their trace, as showcased in Figure˜4 (middle and right column). However, sometimes the situation is more complicated, as showcased in the following example: Consider Figure˜5, which depicts a cycle $C$ on 6 vertices, and a possible realization $r(C)$ with Jordan arcs. Two cells $F,F^{\prime}$ of the arrangement are highlighted. Now, note that the outer cell of the arrangement only touches the Jordan curves corresponding to vertices 1,5,6. This might be considered counter-intuitive, since one could maybe expect the outer boundary curve to touch all Jordan curves for vertices 1 …6. Furthermore, it is not clear what should be the interior bounding curve, since there are two viable candidates, either inside cell $F$ or cell $F^{\prime}$. Since it is not clear what should be the interior or exterior boundary curve in this case, let us simply make the following definition: For any cell $F$ in the arrangement, we define the boundary curve of $F$ to be the curve that closely follows the boundary of $F$ in the arrangement $r(C)$, such that the same properties (i) – (iii) from above are met. Note that since we assumed our Jordan curves to be piecewise differentiable, it is possible to follow the boundary of $F$ close enough such that properties (i) – (iii) are met. Analogously to above, we define the trace of $F$, denoted by $\operatorname{trace}(F)$ as the order of encountered elements that are close to the boundary curve of $F$.
The main insight we offer in this section is that even in the more general case, the trace of $F$ has a very restricted structure. We show in the following lemma that if the cycle $C$ has $n$ vertices, then for any arbitrary cell $F$ of size at least 5 it is always possible to partition the sequence $\operatorname{trace}(F)$ into $n$ pieces, such that the $i$-th piece is either empty or only contains the symbols $i,i+1$ (mod n). In particular, this implies that the order of elements along the trace of $F$ in any geometric realization is fixed, up to possible displacement of either +0/+1 (and up to possible cyclic shift or reflection). Formally, we have the following statement. For notational convenience, all cycle symbols in the lemma are to be understood modulo $n$ and start from 1, that is $n+1=1$ and 11=n1-1=n.
Figure 5: A geometric realization of the cycle on 6 vertices using Jordan arcs.
@@
Let $C\subseteq V(G)$ be an induced cycle of length $n\geq 4$ in the graph $G$, and let $r$ be a geometric realization of $G$ with Jordan curves. Let $F$ be any cell in the arrangement $r(C)$, and let $s:=\operatorname{trace}(F)$ be its trace. If the sequence $s$ contains at least 5 different symbols, then there is a cyclic shift and possible reflection $\tilde{s}$ of $s$, such that we can rewrite $\tilde{s}=s_{1}s_{2}\dots s_{n}$ such that each $s_{i}\in\{i,i+1\}^{\star}$ is nonempty and contains only symbols $i,i+1$.
@@
We fist start by proving a weaker statement. We claim that there exist $n$ sequences $\overline{s}_{1},\dots,\overline{s}_{n}$ such that $\overline{s}_{i}\in\{i-1,i,i+1\}$ and such that $\overline{s}_{1},\dots,\overline{s}_{n}$ together "cover" the cyclic sequence $s$. More precisely, let $k:=|s|$ be the length of the trace, and consider a cycle $D$ on $k$ vertices (not related to the graph $G$). We can consider the cycle $D$ as labeled with the cyclic sequence $s=\operatorname{trace}(F)$. Our goal is to associate to each $i$ a cyclic arc $A_{i}$ of the cycle $D$, such that $|A_{i}|=\overline{s}_{i}$, and such that if we label the arc $A_{i}$ with the sequence $\overline{s}_{i}$ for each $i$, then in all the places where the arcs overlap, the strings agree, and so that all arcs together completely cover $D$ and their labels together result in the sequence $s$. Formally, we make the following claim: It is possible to find $\overline{s}_{1},\dots,\overline{s}_{n}$ such that
  • $\overline{s}_{i}\in\{i-1,i,i+1\}^{\star}$ for all $i=1,\dots,n$.
  • $\overline{s}_{1},\dots,\overline{s}_{n}$ together cover $s$.
  • If $s=\operatorname{trace}(F)$ does not contain any symbol $i$, then $\overline{s}_{i}=\varepsilon$ is the empty sequence
  • If $s=\operatorname{trace}(F)$ contains the symbol $i$, then $\overline{s}_{i}$ contains all occurrences of symbol $i$ and furthermore has symbol $i$ as its first and last letter.
We now prove the claim: Consider for each $i$ all occurrences of symbol $i$ on the cycle. The situation is depicted in Figure˜6(a) for $i=1$. The set of occurrences of symbol $i$ splits the cycle into segments. Consider all occurrences of symbols $X_{i}:=\{1,\dots,n\}\setminus\{i-1,i,i+1\}$ (that is, symbols $3,\dots,n-1$ in Figure˜6(a)). We claim that all these symbols must lie inside at most one segment. Indeed, assume there are $a,a^{\prime}\in X_{i}$ in different segments. Since the subgraph induced by $X_{i}$ is connected, we have that $r(X_{i})$ is a connected subset of the plane, and so $a,a^{\prime}$ are connected by $r(X_{i})$. However, $r(X_{i})$ by definition does not intersect $r(i)$. But this is impossible, since $r(i)$ is a Jordan curve connecting all appearances of symbol $i$, and so $r(i)$ splits all segments from each other, a contradiction. (Note that by definition of $\operatorname{trace}(F)$, no curve can enter the inner of the boundary of FF/the cycle $D$.)
We can hence define $\overline{s}_{i}$ as the cyclic string starting and ending with symbol $i$ containing all segments except possibly the segment containing some symbol from $X_{i}$ (compare Figure˜6(b)). Then all claimed properties of $\overline{s}_{i}$ are true and we have proven the claim.
Next, we argue that for each pair $i\neq j$, it is impossible that $\overline{s}_{i}$ is completely contained in $\overline{s}_{j}$ (assuming $\overline{s}_{i}\neq\varepsilon$). By this, we mean that the corresponding cycle arcs are subsets of each other, i.e. $A_{i}\subseteq A_{j}$. Indeed, note that if this is the case, then $j\in\{i-1,i+1\}$. So for the sake of notation we can w.l.o.g. assume that $i=1$ and $j=2$. Then consider Figure˜6(c). Consider the symbols before and after $\overline{s}_{2}$. Since $\overline{s}_{2}$ starts and ends with a 2, and since $\overline{s}_{1}$ contains all occurrences of 1 and is contained in $\overline{s}_{2}$, both the symbol before and after $\overline{s}_{2}$ must be a 3. We use here that consecutive symbols on the cycle can differ by only one. Consider now $\overline{s}_{3}$. The object $\overline{s}_{3}$ contains all occurrences of $3$, but cannot contain any 1 since $\overline{s}_{3}\in\{2,3,4\}^{\star}$. So $\overline{s}_{3}$ must connect both 3’s, but since $\overline{s}_{1}\neq\varepsilon$ contains at least one 1, the object $\overline{s}_{3}$ must wrap around the other way of the cycle. Then the whole cycle is contained in $\overline{s}_{2}\cup\overline{s}_{3}$. But then $\operatorname{trace}(F)\in\{1,2,3,4\}^{\star}$. This contradicts our assumption that $\operatorname{trace}(F)$ contains at least 5 different symbols.
We use the observation that no $\overline{s}_{j}$ contains any $\overline{s}_{i}$ in order to show the following structural insights: For all $i=1,\dots,n$, $\overline{s}_{i}=\varepsilon$ is actually impossible. To see this, let $\overline{s}_{i}=\varepsilon$ and w.l.o.g. $\overline{s}_{i+1}\neq\varepsilon$. Then symbol $i$ does not appear and $\overline{s}_{i+1}\in\{i+1,i+2\}$. Both the symbol before and after $\overline{s}_{i+1}\in\{i+1,i+2\}$ must be $i+2$, and so either $\overline{s}_{i+1}$ is contained in $\overline{s}_{i+2}$ or $\operatorname{trace}(F)$ has fewer than $5$ distinct symbols, a contradiction. For the same reason, for each $i=1,\dots,n$ it is impossible that the symbols before and after $\overline{s}_{i}$ are both i1i-1; or are both $i+1$.
Finally, we have arrived at the following conclusion: The strings $\overline{s}_{1},\dots,\overline{s}_{n}$ must appear in exactly this order or its reflection along the cycle $D$. (By the order of appearance, we mean the order of the first time we encounter some element from $\overline{s}_{i}$ when going around the cycle). This follows since the two symbols before and after $\overline{s}_{i}$ are exactly $\{i-1,i+1\}$, and since $\overline{s}_{i}$ contains all symbols $i$.
Finally, we show how to construct $s_{1},\dots,s_{n}$ given $\overline{s}_{1}$. We simply define $s_{i}:=\overline{s}_{i}\setminus\overline{s}_{i-1}$. Then since $\overline{s}_{i-1}$ contains all occurrences of i1i-1, we have that $s_{i}\in\{i,i+1\}$. As argued above, the strings $\overline{s}_{1},\dots,\overline{s}_{n}$ are all nonempty, do not contain each other, and appear in exactly this order along the cycle. Then, we conclude that there is a cyclic shift and reflection $\tilde{s}$ of $s$ such that $s=s_{1}s_{2}\dots s_{n}$. The situation is depicted in Figure˜6(d).
(a) $a,a^{\prime}$ cannot be in different segments.
(b) Definition of $\overline{s}_{i}$ for $i=1$.
(c) $A_{1}\subseteq A_{2}$ is impossible
(d) Definition of $s_{i}$ given $\overline{s}_{1},\dots,\overline{s}_{n}$.
Figure 6: Arguments used in the proof of Lemma˜3
The previous lemma will prove to be immensely helpful, since it has the following implication: If we have an induced cycle $C$, and we select a subset $Y$ of elements of pairwise distance at least two in $C$, then any cell in $r(C)$ that touches at least 5 elements must contain the elements of $Y$ in the same unique order along its trace (up to reflection and rotation). Next, we argue that for our graph $G$ the realizations of certain vertices of the graph must be contained in some cell.